# A Style is Worth One Code: Unlocking Code-to-Style Image Generation with Discrete Style Space

### v1

@@ Abstract
Innovative visual stylization is a cornerstone of artistic creation, yet generating novel and consistent visual styles remains a significant challenge. Existing generative approaches typically rely on lengthy textual prompts, reference images, or parameter-efficient fine-tuning to guide style-aware image generation, but often struggle with style consistency, limited creativity, and complex style representations. In this paper, we affirm that a style is worth one numerical code by introducing the novel task, code-to-style image generation, which produces images with novel, consistent visual styles conditioned solely on a numerical style code. To date, this field has only been primarily explored by the industry (e.g., Midjourney), with no open-source research from the academic community. To fill this gap, we propose CoTyle, the first open-source method for this task. Specifically, we first train a discrete style codebook from a collection of images to extract style embeddings. These embeddings serve as conditions for a text-to-image diffusion model (T2I-DM) to generate stylistic images. Subsequently, we train an autoregressive style generator on the discrete style embeddings to model their distribution, allowing the synthesis of novel style embeddings. During inference, a numerical style code is mapped to a unique style embedding by the style generator, and this embedding guides the T2I-DM to generate images in the corresponding style. Unlike existing methods, our method offers unparalleled simplicity and diversity, unlocking a vast space of reproducible styles from minimal input. Extensive experiments validate that CoTyle effectively turns a numerical code into a style controller, demonstrating a style is worth one code.
@@ 1 Introduction
Innovative visual stylization is a cornerstone of artistic creation, enabling the expression of unique identities and emotions across digital art, design, and media. With the advent of generative models, particularly diffusion models [flux2024, labs2025flux, rombach2022latentdiffusion, huang2024dialoggen, saharia2022photorealistic, podell2023sdxl], users can generate high-fidelity images.
Accordingly, controlling generative models to produce images with a specific style has emerged as a critical challenge attracting growing research interest.
Figure 1: Unlike previous methods, CoTyle uses a numerical style code to represent a style, eliminating the need for complex prompts, images, or LoRAs, and allowing easy creation of unique styles just modifying the code. “Creativity”, “Consistency”, and “Transferability” refer to a model’s ability to (1) generate entirely novel styles, (2) produce multiple images in the same style reliably, and (3) apply styles using simple, user-friendly representations.
As shown in Fig. 1, naively using only textual prompts to guide text-to-image diffusion models (T2I-DMs) in generating images with specific styles often fails to achieve the desired results and does not ensure consistency across different generated images. Consequently, some studies [labs2025flux, wu2025uso, wu2021stylealign, xu2025stylessp, wang2024instantstyle, susladkar2025d2styler] inject style information into T2I-DMs by using visual style reference images or parameter-efficient fine-tuning (e.g. Low-Rank Adaptation, LoRA [hu2022lora]) containing specific style information [song2025omniconsistency, xing2024csgo, sohn2023styledrop, shah2024ziplora, frenkel2024blora]. However, these methods suffer from two limitations: (1) they rely on existing styles, lacking the ability to generate novel styles not yet present in the world, limiting their creativity. (2) each target style must be specified either via a pixel-level reference image or a complex set of LoRA , which imposes a burden on users and limits practical transferability.
@@
The core concept of CoTyle is to design a style generator capable of producing novel style embedding that guides the T2I-DM in generating images with a specific style. Specifically, we begin by training a discrete style codebook [esser2021vqgan, van2017vqvae, ma2025unitok] with contrastive loss on paired style images, with the objective of mapping images sharing the same style to similar distribution. This process allows the discrete codebook to pool stylistic information from image features while suppressing irrelevant content. The trained codebook is integrated into a pre-trained T2I-DM. We train the T2I-DM to condition its generation on the codebook’s output. This results in the T2I-DM capable of image-conditioned stylization. To unlock code-to-style generation, we train an autoregressive transformer as a style generator. It learns the distribution of style indices derived from the style codebook under a next-token prediction objective, effectively modeling sequences of indices that represent coherent styles. Thus, the autoregressive model serves as an unconditional style generator, producing novel style index sequences. During inference, a random token, which is fixed by input code, initializes the autoregressive generation process. The model predicts a complete sequence of indices. These indices act as conditioning signals to guide the T2I-DM, enabling the generation of images with a consistent, code-defined style.
To evaluate our method, we employ the CSD [somepalli2024csd] to assess both style consistency and diversity. Extensive experiments validate the efficiency of CoTyle, which generates images in target styles without requiring reference images, lengthy prompts, or model fine-tuning. Furthermore, we extend CoTyle to support image-driven stylized image generation and style interpolation. By offering unprecedented simplicity and fine-grained control over style specification, CoTyle unlocks the potential of code-to-style generation.
@@
  • We propose CoTyle, the first open-source framework that achieves code-to-style generation by learning a discrete style codebook and an autoregressive style generator.
@@ 2 Related Work
Conditioned Image Generation. Conditioned image generation produces visual content conditioned on user-provided control signals. Early methods [dalle3, saharia2022imagen, rombach2022latentdiffusion] established text-based conditioning for text-to-image generation, while subsequent work introduced richer signals. ControlNet [zhang2023controlnet] and T2I-Adapter [mou2024t2i-adapter] provide structural guidance via edge/depth maps or sketches, and DreamBooth [ruiz2023dreambooth] with Textual Inversion [gal2022textualinversion] enable identity-specific generation from few-shot examples. Common conditioning mechanisms include noise inversion [cao2025causalctrl, song2020ddim], adapter modules [guo2024pulid, ye2023ipadapter, betterfit], parameter-efficient fine-tuning [frenkel2024blora, shah2024ziplora, liu2025separate, duan2024tuning], and token concatenation [tan2025ominicontrol, tan2025ominicontrol2, liu2025omnidish, boow, zhang2025grouprelativeattentionguidance]. Few studies leverage vision-language models to encode visual information as prompt embeddings; while Qwen-Image [wu2025qwenimage] does so, it is primarily designed for image editing.
@@
3 Task Definition
In this paper, we propose a novel task, code-to-style image generation. Given a numerical style code and a textual prompt, the system generates an image that realizes the specified artistic style while semantically adhering to the content described in the prompt. Existing research on stylistic image generation can be categorized into several mainstream approaches, yet significant discrepancies exist between these methods and the proposed code-to-style generation, as illustrated in Fig. 1. The radar chart in the figure compares representative methods across three key evaluation dimensions: (1) Consistency. Images generated using the same style condition should exhibit high style consistency. (2) Creativity. The model should be capable of generating styles that do not exist in the real world. (3) Users can transmit style information to other users or systems via a concise and portable representation, enabling reproducible generation. A straightforward approach is to generate stylistic images using textual prompts that describe the desired style. Images generated from identical textual style descriptions often exhibit substantial variations in visual appearance, leading to poor stylistic consistency. Another category of methods either leverages a reference image as a style condition to guide the model in generating images with matching styles [wu2025uso, lei2025stylestudio, xing2024csgo, wang2024instantstyle-plus] or inserts a pre-trained style LoRA (Low-Rank Adaptation [hu2022lora]) into the generative model to enable specific style generation [frenkel2024blora, shah2024ziplora]. Nevertheless, both methods inherently rely on existing style images (either as reference images or LoRA training data), rendering them incapable of creating novel, unseen artistic styles. Furthermore, transferring style information for reproducible generation via these methods requires sharing pixel-domain reference images or complex LoRA weights, resulting in poor transferability.
In contrast, the proposed code-to-style generation addresses these limitations: (1) similar to image-conditioned methods, it guarantees high stylistic consistency across multiple images generated with the same style code; (2) it enables the creation of novel artistic styles based on arbitrary style codes, achieving strong creativity; (3) 3he numerical style code serves as a concise and portable representation of the target style, ensuring excellent transferability for seamless sharing and reproducible generation.
@@ 4.3 Code-to-Style Image Generation
To further enhance the intensity and diversity of the style, we propose a sampling strategy. We analyzed the frequency of codebook index across a collection of images, and find that certain indices exhibit significantly higher selection frequencies than others (see Appendix for details). Empirical results (Fig. 5) show that high-frequency indices represent a form of “arbitrary” style information, just like meaningless placeholders. Sampling exclusively from high-frequency indices yields images with no specific style. Therefore, during inference, we down-weight the logits corresponding to these indices by muliplying a suppression coefficient. The suppression coefficient can be formulated as,
Inputs: style code $s$, prompt $y$
Given: style codebook $C$, codebook size $K$, number of tokens NN
@@ 5.1 Implementation Details
During the codebook training phase, we set the codebook vocabulary size to 1,024 and the embedding dimension to 64. The model was trained for 20,000 steps with a batch size of 128 and a learning rate of 1e-5. In the DiT training stage, we initialized the model from a pre-trained T2I-DM and trained it for 60,000 steps with a batch size of 64 and a learning rate of 4e-6. For the style generator, we employed the Qwen2-0.5B [qwen2] architecture but trained it from scratch without loading any pre-trained weights. This component was trained for 100,000 steps with a batch size of 64 and a learning rate of 1e-5. We resized style references to 392×392 during training, as stylistic features are robust to geometric transformations. Each image was then encoded into 196 style tokens, represented by 196 discrete style indices (e.g. $N=196$).
Table 1: Quantitative comparison with other methods. “ * ” denotes CoTyle conditioned on style images. Bold highlights the best score; underlines “    ” indicate the second-highest, omitted in code-to-style evaluation for brevity. CLIP-T measures text-image alignment. For image-conditioned methods, diversity is constrained by the reference image and is neither measurable nor meaningful to evaluate.
@@ 5.2 Benchmark
For code-to-style generation, we randomly sampled 500 codes, generating 4 images per code, resulting in 2,000 images for evaluation. For image-conditioned generation, we constructed an evaluation set of 500 prompt-image pairs.
Following [wu2025uso, xing2024csgo], we adopted CSD [somepalli2024csd] to evaluate these methods. For style consistency (Consistency), we computed the CSD score among images generated from the same code for CoTyle and Midjourney [mj]. For image-conditioned methods [wu2025uso, wang2024instantstyle-plus, xing2024csgo, lei2025stylestudio, qi2024deadiff], we measured the CSD score between their outputs and reference images. To evaluate style diversity (Diversity), we randomly sampled multiple codes, generated corresponding images, and computed pairwise CSD scores among them. We define the Diversity score as 1 minus the CSD score, so that higher values indicate greater diversity. The Diversity score measures the degree of stylistic diversity across styles corresponding to different codes. Therefore, this metric is specifically designed for the code-to-style task. Additionally, we evaluate text-image alignment using CLIP text-image similarity (CLIP-T) [wu2025uso, xing2024csgo, lei2025stylestudio] and assess the aesthetic quality of generated images via the recently popular QualityCLIP [agnolucci2024quality].
@@ 5.4 Results
Figure 4: Qualitative comparison. CoTyle is not only capable of generation conditioned on style codes but also supports style images. The text highlighted in red underscores that our model can faithfully follow the input text while simultaneously generating the specified style.
Code-to-style image generation. We presented a comparison between CoTyle and the closed-source Midjourney [mj] on the code-to-style task in Fig. 3 and Tab. 1, demonstrating that our method demonstrated significantly superior style consistency. CoTyle achieved better style consistency, where different images generated from the same code exhibited highly similar styles. However, CoTyle showed slightly inferior diversity, which may be attributed to the suboptimal training of the style generator. This represented a potential direction for future improvement.
@@ 5.5 Ablation
Effect of style injection through textual branch. CoTyle leverages a VLM as the text encoder to inject style information through the textual branch. For comparison, we implement style injection through the visual branch using the OmniControl [tan2025ominicontrol, tan2025ominicontrol2], i.e., by concatenating noise features with style condition features along the token dimension. The quantitative results in Tab. 2 confirm this perspective. More visual analysis can be found in Appendix.
@@
Effect of style loss. During style codebook training, the style loss comprised a contrastive loss for style learning and a reconstruction loss. We validated the necessity of both the contrastive and reconstruction components, The contrastive loss helped the model better extract style information, enhancing style consistency, while the reconstruction loss aligns the style embedding with the original ViT feature distribution, improving overall performance. The results in Tab. 3 support this hypothesis.
Effect of frequency suppression. We found that sampling exclusively from high-frequency indices resulted in images devoid of any style information, as illustrated in Fig. 5. We hypothesized that high-frequency indices might function as meaningless placeholders. Therefore, we propose high-frequency suppression strategy. As shown in Tab. 4, without this strategy, the generated images exhibit low style diversity, as many style codes tend to produce photorealistic results.
@@ 6 Conclusion
In this paper, we introduce code-to-style image generation, a novel paradigm that enables controllable generation of diverse visual styles through numerical style codes. We propose CoTyle, the first open-source framework to support code-to-style generation. CoTyle employs a dedicated style generator trained to produce novel and semantically meaningful style representations from numerical style codes. Then, a trained diffusion model generates images with specific style conditioned on the style representation. Further, we extend CoTyle to support additional functionalities, including image-conditioned style generation and style interpolation fusion. Extensive experiments demonstrate that CoTyle effectively maps discrete codes to diverse visual styles. Our work establishes a new paradigm and opens up future research in discrete stylistic representations.
@@ 8 Analysis of High-frequency Index
9 Analysis of Style Injection
As noted in the main paper, we propose injecting style information via the textual branch. Specifically, we employ a vision language model (VLM) as the text encoder to encode style reference images. While quantitative results are presented in the experiments of the main paper, we further provide qualitative results here to complement the analysis (see Fig. A2). In the figure, the first column shows style reference images, conditioning inputs for the text-to-image diffusion model; the second column presents results of style injection via the visual branch, following a scheme similar to OminiControl [tan2025ominicontrol]; the third column depicts CoTyle’s outputs, which inject style information through the textual branch using a VLM [qwen2.5-VL] as the text encoder. As shown in the first row of Fig. A2, visual branch injection captures the red tones of the paper-cut style but fails to recognize semantic-style elements (e.g., circular contours). For the second row, visual branch injection only captures warm-toned cues, whereas textual branch injection enables the model to correctly generate a human figure composed of crystal blocks.
Figure A2: We compare injecting style through textual branch with the existing method through visual branch. Injecting style from the textual branch better preserves semantic information.

### v3

@@ Abstract
Innovative visual stylization is a cornerstone of artistic creation, yet generating novel and consistent visual styles remains a significant challenge. Existing generative approaches typically rely on lengthy textual prompts, reference images, or parameter-efficient fine-tuning to guide style-aware image generation, but often struggle with style consistency, limited creativity, and complex style representations. In this paper, we consider the code-to-style image generation task, which aims to produce images with novel and consistent visual styles specified by only a numerical code. To date, this field has only been primarily explored by the industry (e.g., Midjourney), with no open-source research from the academic community. To fill this gap, we propose CoTyle, the first open-source method for this task. Specifically, we first train a discrete style codebook from a collection of images to extract style embeddings. These embeddings serve as conditions for a text-to-image diffusion model (T2I-DM) to generate stylistic images. Subsequently, we train an autoregressive style generator on the discrete style embeddings to model their distribution, allowing the synthesis of novel style embeddings. During inference, a numerical style code is mapped to a unique style embedding by the style generator, and this embedding guides the T2I-DM to generate images in the corresponding style. Extensive experiments validate that CoTyle effectively converts a numerical code into a style controller, demonstrating a style is worth one code. Compared to existing methods, the stylized images generated by our method are more diverse and consistent, unlocking a vast space of reproducible styles from minimal input.
@@ 1 Introduction
Innovative visual stylization is a cornerstone of artistic creation, enabling the expression of unique identities and emotions across digital art, design, and media. With generative models, e.g., diffusion models [flux2024, labs2025flux, rombach2022latentdiffusion, huang2024dialoggen, saharia2022photorealistic, podell2023sdxl], users can generate high-fidelity images with arbitrary styles.
Existing works on stylistic image generation can be categorized into several groups according to the style specification type, including methods specifying styles by textural prompt, visual image, Low-Rank Adaptation (LoRA) [hu2022lora], as illustrated in Fig. 1. The radar chart in the figure compares representative methods across three key evaluation dimensions, including consistency, creativity and reproducibility. Methods specified style by textural prompts usually exhibit poor stylistic consistency, as images generated from identical textual style descriptions often exhibit substantial variations in visual appearance. Other methods either leverage the style of reference image as the style definition  [wu2025uso, lei2025stylestudio, xing2024csgo, wang2024instantstyle-plus] or insert a pre-trained style LoRA [hu2022lora] into the generative model to enable specific style generation [frenkel2024blora, shah2024ziplora]. Nevertheless, both methods rely on existing style images (either as reference images or LoRA training data), rendering them incapable of creating novel, unseen artistic styles, i.e. they own poor creativity. Furthermore, transferring style information for reproducible generation via these methods requires sharing pixel-level reference images or complex LoRA weights, resulting in poor reproducibility. Thus, existing methods cannot simultaneously achieve high consistency, creativity and reproducibility.
Figure 1: Different to previous methods, CoTyle uses a numerical style code to represent a style, eliminating the need for complex prompts, images, or LoRAs, and allowing easy creation of unique styles just modifying the code. “Creativity”, “Consistency”, and “Reproducibility” refer to a model’s ability to (1) generate novel styles, (2) produce multiple images in the same style consistently, and (3) reproduce styles using simple, user-friendly style definitions.
@@
The core concept of CoTyle is to design a style generator capable of producing novel style embedding that guides the T2I-DM in generating images with a specific style. Specifically, we begin by training a discrete style codebook [esser2021vqgan, van2017vqvae, ma2025unitok] with contrastive loss on paired style images, with the objective of mapping images sharing the same style to similar distribution, while pushing apart the distributions of distinct styles. This process allows the discrete codebook to pool stylistic information from image features while suppressing irrelevant content. The trained codebook is integrated into a pre-trained T2I-DM. We train the T2I-DM to condition its generation on the codebook’s output. This results in the T2I-DM capable of image-conditioned stylization. To unlock code-to-style generation, we train an autoregressive transformer as a style generator. It learns the distribution of style indices derived from the style codebook under a next-token prediction objective, effectively modeling sequences of indices that represent coherent styles. Thus, the autoregressive model serves as an unconditional style generator, producing novel style index sequences. During inference, a random token, which is fixed by input code, initializes the autoregressive generation process. The model predicts a complete sequence of indices. These indices act as conditioning signals to guide the T2I-DM, enabling the generation of images with a consistent, code-defined style.
To evaluate our method, we employ the CSD [somepalli2024csd] to assess both style consistency and diversity. Extensive experiments validate the efficiency of CoTyle, which generates images in target styles without requiring reference images, lengthy prompts, or model fine-tuning. Furthermore, we extend CoTyle to support image-driven stylized image generation and style interpolation. As illustrated in Fig. LABEL:fig:teaser, CoTyle addresses limitations existing in previous works: 1) similar to image-conditioned methods, it guarantees high stylistic consistency across multiple images generated with the same style code; 2) it enables the creation of novel artistic styles based on arbitrary style codes, achieving strong creativity; 3) the numerical style code serves as a concise and portable representation of the target style, ensuring excellent reproducibility for seamless sharing and reproducible generation.
@@
  • We propose CoTyle, the first open-source framework that achieves code-to-style generation by learning a discrete style codebook and an autoregressive style generator. Further, we expand it to facilitate image-conditioned generation and supports style interpolation.
@@ 2 Related Work
Conditioned Image Generation. Conditioned image generation produces visual content conditioned on user-provided control signals. Early methods [dalle3, saharia2022imagen, rombach2022latentdiffusion] established text-based conditioning for text-to-image generation, while subsequent work introduced richer signals. ControlNet [zhang2023controlnet] and T2I-Adapter [mou2024t2i-adapter] provide structural guidance via edge/depth maps or sketches, and DreamBooth [ruiz2023dreambooth] with Textual Inversion [gal2022textualinversion] enable identity-specific generation from few-shot examples. Common conditioning mechanisms include noise inversion [cao2025causalctrl, song2020ddim], adapter modules [guo2024pulid, ye2023ipadapter, betterfit], parameter-efficient fine-tuning [frenkel2024blora, shah2024ziplora, liu2025separate], and token concatenation [tan2025ominicontrol, tan2025ominicontrol2, liu2025omnidish, boow, zhang2025grouprelativeattentionguidance]. Few studies leverage vision-language models to encode visual information as prompt embeddings; while Qwen-Image [wu2025qwenimage] does so, it is primarily designed for image editing.
@@ 3.3 Code-to-Style Image Generation
To further enhance the intensity and diversity of the style, we propose a sampling strategy. We analyzed the frequency of codebook index across a collection of images, and find that certain indices exhibit significantly higher selection frequencies than others (see Appendix for details). Empirical results (Fig. 6) show that high-frequency indices represent a form of “arbitrary” style information, just like meaningless placeholders. Sampling exclusively from high-frequency indices yields images with no specific style. Therefore, during inference, we down-weight the logits corresponding to these indices by multiplying a suppression coefficient. The suppression coefficient can be formulated as,
Inputs: style code $n$, prompt $y$
Given: style codebook $C$, codebook size $K$, number of tokens NN, suppression coefficient $s(i)$
@@ 4.1 Implementation Details
During the codebook training phase, we set the codebook vocabulary size to 1,024 and the embedding dimension to 64. The model was trained for 20,000 steps with a batch size of 128 and a learning rate of 1e-5. In the DiT training stage, we initialized the model from pretrained Qwen-Image [wu2025qwenimage] and trained it for 60,000 steps with a batch size of 64 and a learning rate of 4e-6. For the style generator, we employed the Qwen2-0.5B [qwen2] architecture but trained it from scratch without loading any pre-trained weights. This component was trained for 100,000 steps with a batch size of 64 and a learning rate of 1e-5. We resized style references to 392×392 during training, as stylistic features are robust to geometric transformations. Each image was then encoded into 196 style tokens, represented by 196 discrete style indices (i.e. $N=196$).
Table 1: Quantitative comparison with other methods. “ * ” denotes CoTyle conditioned on reference images. Bold highlights the best score; underlines “    ” indicate the second-highest, omitted in code-to-style evaluation for brevity. CLIP-T measures text-image alignment. For image-conditioned methods, diversity is constrained by the reference image and is neither measurable nor meaningful to evaluate.
@@ 4.2 Benchmark
For code-to-style generation, we randomly sampled 500 codes, generating 4 images per code, resulting in 2,000 images for evaluation. For image-conditioned generation, we constructed an evaluation set of 500 prompt and reference image pairs.
Following [wu2025uso, xing2024csgo], we adopted CSD [somepalli2024csd] to evaluate image-conditioned methods. For style consistency (Consistency), we computed the CSD score among images generated from the same code for CoTyle and Midjourney [mj]. For image-conditioned methods [wu2025uso, wang2024instantstyle-plus, xing2024csgo, lei2025stylestudio, qi2024deadiff], we measured the CSD score between their outputs and reference images. To evaluate style diversity (Diversity), we randomly sampled multiple codes, generated corresponding images, and computed pairwise CSD scores among them. We define the Diversity score as 1 minus the CSD score, so that higher values indicate greater diversity. The Diversity score measures the degree of stylistic diversity across styles corresponding to different codes. Therefore, this metric is specifically designed for the code-to-style task. Additionally, we evaluate text-image alignment using CLIP text-image similarity (CLIP-T) [wu2025uso, xing2024csgo, lei2025stylestudio] and assess the aesthetic quality of generated images via the recently popular QualityCLIP (Aesthetics) [agnolucci2024quality].
@@ 4.4 Results
Code-to-style image generation. We presented a comparison between CoTyle and the closed-source Midjourney [mj] on the code-to-style task in Fig. 3 and Tab. 1, demonstrating that our method achieves significantly superior style consistency. CoTyle achieved better style consistency, where different images generated from the same code exhibited highly similar styles. However, CoTyle showed slightly inferior diversity, which may be attributed to the insufficient breadth of the training dataset, constraining CoTyle’s ability to learn a wider spectrum of stylistic variations. This represented a potential direction for future improvement.
@@
Figure 4: We compare injecting style through textual branch with the existing method through visual branch. Injecting style from the textual branch better preserves semantic information.
@@ 4.5 Ablation
Effect of style injection through textual branch. CoTyle leverages a VLM as the text encoder to inject style information through the textual branch. For comparison, we implement style injection through the visual branch using the OminiControl [tan2025ominicontrol, tan2025ominicontrol2], i.e., by concatenating noise features with style condition features along the token dimension. The quantitative results in Tab. 2 confirm this perspective. As shown in the first row of Fig. 4, visual branch injection captures the red tones of the paper-cut style but fails to recognize semantic-style elements (e.g., circular contours). For the second row, visual branch injection only captures warm-toned cues, whereas textual branch injection enables the model to correctly generate a human figure composed of crystal blocks.
@@
Figure 5: Qualitative comparison. CoTyle is not only capable of generation conditioned on style codes but also supports style images. Our model can faithfully follow the input text while simultaneously generating the specified style.
@@
Effect of style loss. During style codebook training, the style loss $L_{style}$ comprised a contrastive loss for style learning and a reconstruction loss. We validated the necessity of both the contrastive and reconstruction components, The contrastive loss helped the model better extract style information, enhancing style consistency, while the reconstruction loss aligns the style embedding with the original ViT feature distribution, improving overall performance. The results in Tab. 3 support this hypothesis.
Effect of frequency suppression. We found that sampling exclusively from high-frequency indices resulted in images devoid of stylistic diversity, as shown in Fig. 6. We hypothesized that high-frequency indices might function as meaningless placeholders. Therefore, we propose high-frequency suppression strategy. Tab. 4 illustrates that without this strategy, the generated images exhibit low style diversity, as most style codes tend to produce photorealistic results. More analyses are provided in the Appendix.
@@ 5 Conclusion
In this paper, we introduce code-to-style image generation, a novel task that enables stylistic image generation through numerical style codes. We propose CoTyle, the first open-source framework to support this task. CoTyle employs a dedicated style generator trained to produce novel style embedding from numerical style codes. Then, a diffusion model generates images with specific style conditioned on the style embedding. Further, we extend CoTyle to support additional functionalities, including image-conditioned style generation and style interpolation. Extensive experiments demonstrate that CoTyle effectively maps discrete codes to diverse visual styles. Our work establishes opens up future research in discrete stylistic representations.
@@ 7 Analysis of High-frequency Index
8 Implications and Insights
Figure A2: The impact of token selection strategy on the generation results during style interpolation.
In this paper, we not only propose a novel approach to code-to-style generation but also offer several insightful observations worth further discussion.
First, we introduce a discrete style codebook to extract and represent stylistic information from images in a discrete latent space. Building upon this representation, we demonstrate that smooth style interpolation can be achieved by blending different style codes. As shown in Fig. LABEL:, we observe that the specific token selection strategy during interpolation has minimal impact on the generated results, regardless of whether it is achieved through random sampling or splitting tokens evenly between the two styles (e.g., first half from style A and second half from style B). This indirectly highlights a key distinction between our learned style representation and traditional image representations: the stylistic information is invariant to the order of its tokens. This property aligns naturally with human intuition about style, which is typically perceived as a holistic attribute rather than a sequence-dependent structure. Style interpolation is just one emergent application enabled by our learned style representation; the structural properties of this discrete space may facilitate the development of additional, more sophisticated, and practically valuable functionalities. Furthermore, this discrete feature extraction paradigm holds promise for extension to other modalities, such as audio.
Second, we advocate for injecting style information from the textual modality rather than the visual. We argue that human perception of style is inherently semantic rather than purely chromatic; thus, conditioning on text enables the model to capture stylistic attributes that better align with human intuition.
Finally, code-to-style image generation is a task of substantial practical potential, yet it has remained unexplored in the academic literature. We present the first comprehensive framework to address this problem and publicly release both the model weights and source code, which we hope will inspire further research in this emerging direction.
9 Limitations
The CoTyle framework represents a significant advancement in the domain of code-to-style image generation. However, several limitations should be acknowledged and explored in future research.
Firstly, while the experimental results indicate that the stylistic diversity of images generated by CoTyle is commendable, it remains somewhat lower than that observed in Midjourney. This outcome is likely linked to the training regimen of the autoregressive style generator, which relies on a discrete style codebook. This codebook, constructed from a limited dataset of curated style-paired images, may not fully encapsulate the extensive variability inherent in human artistic expressions. Expanding the diversity of training data with a broader range of abstract and diverse artistic sources, or implementing more advanced generative architectures like mixture-of-experts models, may enhance the representation of complex, multi-modal style distributions.
Secondly, the resolution and creativity of the generated styles are largely dictated by the quality of the style codebook. Despite utilizing contrastive loss to differentiate styles, the necessity to quantize these into a finite codebook inevitably entails some loss of detailed stylistic subtleties. As a result, certain intricate or hybrid nuances found in reference images may be attenuated or omitted, potentially affecting the expressiveness of the generated outputs.
Finally, although the numerical style code offers a robust framework for reproducibility, the exploration of aesthetically or artistically engaging styles presently involves a stochastic search across a vast code space. Introducing more intuitive mechanisms for navigating or modifying this latent style space could significantly enhance both user agency and practical applicability.
Although CoTyle still has some areas for improvement, as the first open-source work on the task of code-to-style generation, it introduces an overall framework for addressing this task, holding the potential to inspire subsequent research in the field of code-to-style generation.
10 More Results
10.1 Style Interpolation
10.2 Code-to-Style Generation